# references.bib RefAV: Towards Planning-Centric Scenario Mining

### v2

@@ Appendix B Scenario Mining Annotation Tool
In order to manually annotate interesting scenarios, we build a custom web app with Claude 3.7 (cf. Fig 5). To construct a scenario, the user selects an object by clicking on a point within the image frame, writes a natural language description, and selects the start and end frames of the referred objects correspond to the prompt. We project a ray originating from the 2D point and find the ground truth 3D bounding box centroid closest to the ray. This tool allows us to quickly generate multi-object and multi-camera referential tracks.
Appendix C Referential Tracking Evaluation of Related and Other Objects
@@
Appendix D Impact of Sampling Rate on RefProg
@@
Appendix E More Dataset Statistics
@@
| Positive Prompts | 7910 |
| Negative Prompts | 2090 |
@@
| Only Ego is Referred | 827 |
| Only Ego is Related | 1200 |
Appendix F Summary of CVPR 2025 Competition Top Performers
We summarize the contributions of top teams below. We present full technical reports and code here.
Zeekr UMCV extends RefProg with global context-aware generation to processes all queries collectively for improved reasoning and consistency. They also propose a multi-agent refinement system where a secondary refiner agent iteratively debugs and enhances the generated code.
Mi3 UCM proposes two key improvements to RefProg: (1) Fault-Tolerant Iterative Code Generation (FT-ICG), which refines faulty code by re-prompting the LLM with error feedback until successful execution, and (2) Enhanced Prompting for Spatial Relational Functions (EP-SRF), which clarifies function semantics to prevent misinterpretations of spatial relationships between objects.
ZXH refines RefProg by prompting LLMs with a batch of queries and in-context examples instead of prompting for each query one at a time. Importantly, ZXH’s approach relies on in-context learning and batch prompting to improve consistency and efficiency.
Appendix G API Listing

### v3

@@ Appendix D Scenario Mining Annotation Tool
In order to manually annotate interesting scenarios, we build a custom web app with Claude 3.7 (Figure 5). To construct a scenario, the user selects an object by clicking on a point within the image frame, writes a natural language description, and selects the start and end frames of the referred objects correspond to the prompt. We project a ray originating from the 2D point and find the ground truth 3D bounding box centroid closest to the ray. This tool allows us to quickly generate multi-object and multi-camera referential tracks.
Appendix E Referential Tracking Evaluation of Related and Other Objects
@@
Appendix F Impact of Sampling Rate on RefProg
@@
Appendix G More Dataset Statistics
Figure 6 quantifies important characteristics of RefAV such as the number of unique scenario descriptions, the number of positive and negative examples, the number of procedurally and manually defined scenarios, the number of scenarios that include object relationships, and the number of scenarios directly involving the ego-vehicle. We use the catch-all term “expression” for scenario description to better compare RefAV to other datasets. A positive match occurs when the data log contains at least one referred object corresponding to the scenario description. Figure 7 shows the number of positive and negative examples for representative scenario descriptions. In Figure 8, we evaluate the spatial distribution of referred objects.
@@ Appendix G More Dataset Statistics
| Unique | 407 |
| Positive Match | 7910 |
| Negative Match | 2090 |
@@
| Ego is the only Referred Object | 827 |
| Ego is the only Related Object | 1200 |
Appendix H Summary of Competition Top Performers
We summarize the contributions of top teams below.
Team Zeekr UMCV extends RefProg with global context-aware generation to processes all queries collectively for improved reasoning and consistency. They also propose a multi-agent refinement system where a secondary refiner agent iteratively debugs and enhances the generated code.
Team Mi3 UCM proposes two key improvements to RefProg: (1) Fault-Tolerant Iterative Code Generation (FT-ICG), which refines faulty code by re-prompting the LLM with error feedback until successful execution, and (2) Enhanced Prompting for Spatial Relational Functions (EP-SRF), which clarifies function semantics to prevent misinterpretations of spatial relationships between objects.
Team ZXH refines RefProg by prompting LLMs with a batch of queries and in-context examples instead of prompting for each query one at a time. Importantly, this team’s approach relies on in-context learning and batch prompting to improve consistency and efficiency.
Appendix I Examples of RefProg’s Synthesized Programs
We provide examples of RefProg’s correctly synthesized programs below.
Prompt: Pedestrian crossing between stopped buses
Prompt: Group of at least 3 moving bicyclists with 5m from each other
Prompt: Accelerating vehicle changing lanes to the right
We provide examples of RefProg’s incorrectly synthesized programs below.
Prompt: bus in front of ego vehicle obstructing view of red traffic light
The following program fails because the pre-defined atomic functions lack the expressivity to synthesize a correct program. Interestingly, the LLM realizes that the provided functions don’t allow us to determine if a traffic light is being obstructed.
Prompt: bicycle following closely behind a vehicle
The following program fails due to imprecise semantics because the bicycle is not necessarily following the vehicle.
Prompt: vehicle starting to switch lanes but then going back
The following program uses incorrect logic. It fails to combine lane changes and acceleration and uses AND on exclusive sets.
Appendix J API Listing